Propane (C3H8) has a normal boiling point of -42.0°C and a heat of Vaporization of 19.04Kj/mol. What is the Vapour Pressure of propane at 25.0°C?
1
Expert's answer
2020-08-14T07:25:51-0400

T1 = -42.0 + 273.15 = 231.15 K

T2 = 25 + 273.15 = 298.15

∆Hvap = 19.04 kJ/mol = 19040 J/mol

At normal boiling point, the vapor pressure is equal to 1 atm.

p1 = 1 atm

Clausius Clapeyron equation:

ln(p2/p1) = (∆Hvap/R)×(1/T1 – 1/T2)

R — the gas constant (8.3145 J/mol K).

ln(p2/1) = (19040/8.3145)×(1/231.15 – 1/298.15)

ln(p2/1) = 2289.97×(0.00432 – 0.00335)

ln(p2/1) = 2289.97×0.00097

ln(p2/1) = 2.22

p2 = e2.22 = 9.207 atm

Answer: 9.207 atm